# LLM as GNN: Graph Vocabulary Learning for Text-Attributed Graph Foundation Models

### v1

@@ Abstract
Text-Attributed Graphs (TAGs), where each node is associated with text descriptions, are ubiquitous in real-world scenarios. They typically exhibit distinctive structure and domain-specific knowledge, motivating the development of a Graph Foundation Model (GFM) that generalizes across diverse graphs and tasks. Despite large efforts to integrate Large Language Models (LLMs) and Graph Neural Networks (GNNs) for TAGs, existing approaches suffer from decoupled architectures with two-stage alignment, limiting their synergistic potential. Even worse, existing methods assign out-of-vocabulary (OOV) tokens to graph nodes, leading to graph-specific semantics, token explosion, and incompatibility with task-oriented prompt templates, which hinders cross-graph and cross-task transferability. To address these challenges, we propose PromptGFM, a versatile GFM for TAGs grounded in graph vocabulary learning. PromptGFM comprises two key components: (1) Graph Understanding Module, which explicitly prompts LLMs to replicate the finest GNN workflow within the text space, facilitating seamless GNN-LLM integration and elegant graph-text alignment; (2) Graph Inference Module, which establishes a language-based graph vocabulary ensuring expressiveness, transferability, and scalability, enabling readable instructions for LLM fine-tuning. Extensive experiments demonstrate our superiority and transferability across diverse graphs and tasks. The code is available at this URL 111 https://github.com/agiresearch/PromptGFM.
LLM as GNN: Graph Vocabulary Learning for Text-Attributed
Graph Foundation Models
Xi Zhu1thanks: Equal contribution., Haochen Xue211footnotemark: 1, Ziwei Zhao3, Wujiang Xu1, Jingyuan Huang1, Minghao Guo1, Qifan Wang4, Kaixiong Zhou5, Yongfeng Zhang1thanks: Corresponding email: yongfeng.zhang@rutgers.edu. 1 Rutgers University 2 University of Liverpool 3 University of Science and Technology of China 4 Meta AI 5 North Carolina State University
@@ 1 Introduction
Graphs, characterized by their non-Euclidean structures and rich domain-specific knowledge, serve as fundamental representations of complex relational data. Many of them integrate textual information at the node level, forming Text-Attributed Graphs (TAGs), such as citation networks  (Eto, 2019; Hu et al., 2020; Buneman et al., 2021), social networks  (Kempe et al., 2003; Myers et al., 2014), and molecular graphs  (Wieder et al., 2020; Jin et al., 2024a). However, existing solutions heavily depend on task- or dataset-specific training and deployment, limiting their transferability. To address this, we aim to build a Graph Foundation Model (GFM) capable of generalizing across different graphs and tasks for TAGs  (Mao et al., 2024; Xia et al., 2024).
@@
In TAGs, Graph Neural Networks (GNNs) and Large Language Models (LLMs) are employed to handle different modalities, as illustrated in Figure 1. (a) GNN for LLM. GNNs generate structure-aware node embeddings that complement original textual embeddings, improving LLM inference  (Tang et al., 2024; Chai et al., 2023; Liu et al., 2024b). (b) LLM for GNN. LLMs extract additional semantic features or labels from textual data, serving as supervision signals for GNN training  (Chen et al., 2024c; Liu et al., 2024a; Zhu et al., 2024). However, current loosely coupled architectures with two-stage alignment struggle to fully exploit the synergy between GNNs and LLMs, resulting in suboptimal graph-text alignment in TAGs.
@@
Recently, a noteworthy trend has emerged toward implementing LLM as GNN, where graph verbalizers convert graph data into code-like or heuristic prompts, enabling LLMs prompts to understand graph semantics and structure  (Ye et al., 2024; Wang et al., 2024a; Chen et al., 2024a). However, we argue that currently there are no true examples of this category, as they lack the essence of a genuine GNN: the message-passing paradigm. As shown in Figure 2(a), a traditional GNN layer includes neighbor sampling, aggregation-update, and optimization  (Kipf and Welling, 2017; Velickovic et al., 2017). By stacking multiple layers, structure-less embeddings are gradually transformed into structure-rich embeddings with higher-order signals. Due to the absence of these key components, an urgent challenge arises: Can we leverage LLMs to faithfully replicate GNNs to capture both graph semantics and structures simultaneously?
@@
Meanwhile, existing works intuitively treat each graph node as an out-of-vocabulary (OOV) token, leading to graph-specific semantics and uncontrolled token explosion  (Tang et al., 2024; Ye et al., 2024). Worse still, due to vocabulary mismatches, ID-based node embeddings within graphs and language-based token embeddings from templates reside in different feature spaces, resulting in semantic misalignment in LLM inference. This incompatibility further hinders the transferability and scalability of graph-specific knowledge across other graphs and tasks. To enable knowledge transfer, a critical challenge emerges: Can we replace OOV tokens with compatible and universal node representations to build a versatile GFM?
A versatile GFM should be grounded in graph vocabulary learning  (Mao et al., 2024; Cai, 2024), ensuring the following properties: (1) Expressiveness: it encapsulates both semantic and structural information across diverse graphs. (2) Transferability: Every node in any graph should be representable using one or more fundamental units within the vocabulary. (3) Scalability: the vocabulary should be inclusive to accommodate unseen nodes, even those beyond existing graphs. Since natural language is a highly expressive medium made up of meaningful and transferable tokens  (Raffel et al., 2020; Radford et al., 2021; Palo et al., 2023; Wang et al., 2024c), we are inspired to establish a universal graph vocabulary within the text space for node representations. Thus, we propose PromptGFM as a GFM for TAGs as follows.
@@
Graph Understanding Module. To function LLMs as GNNs, we initialize node features with textual attributes and prompt LLMs to explicitly replicate the fine-grained GNN workflow within the text space. As illustrated in Figure 2, we represent graph structure through one-hop neighbor descriptions and design prompts to guide a flexible aggregation-update mechanism. To optimize, we incorporate heuristic prompts to reflect contrastive loss in each layer. Finally, iterative LLM calls simulate message passing, progressively refining verbose textual features into concise yet meaningful textual representations rather than numerical embeddings. Aligning with embedding-based GNN, our prompt-based GNN successfully preserves node semantics while capturing higher-order connections. Consequently, LLMs can operate as GNNs, and GNNs can be interpreted as LLMs, unlocking the full potential of GNN-LLM integration and empowering elegant graph-text alignment.
Graph Inference Module. Having captured semantic and structural information through prompt-based GNNs, we decouple textual representations to establish a graph vocabulary, where each node is mapped to a finite sequence of language-based tokens, essentially as language-based IDs. This vocabulary is universally transferable and scalable across graphs, which resolves semantic irrelevance and graph-text incompatibility. Afterward, these language-based IDs can be incorporated to generate readable and coherent pure-language prompts. Within a multi-instruction fine-tuning framework, we collect diverse instructions across graphs and tasks to effectively fine-tune an LLM, enabling cross-graph and cross-task knowledge transfer. In conclusion, this graph vocabulary eliminates incompatibility and paves the way for general GFMs.
@@ 2 Related Works
GNN-LLM Integration. LLMs have unlocked unprecedented potential for graph machine learning, driving efforts to integrate GNNs and LLMs for modeling TAGs. (a) GNN for LLM. As structure tokenizers, GNNs and graph transformers embed graph topology into textual representations, enhancing semantical understanding in LLMs (Tang et al., 2024; Chai et al., 2023; Liu et al., 2024b). However, effectively coordinating GNN-LLM architectures and co-training remains a challenge. (b) LLM for GNN. LLMs assist GNNs by generating node-level and edge-level labels to address data sparsity issues  (Chen et al., 2024c; Guo et al., 2024; Xia et al., 2024; Shu et al., 2024). Meanwhile, the other line directs LLMs to produce additional features or explanations to overcome semantic deficiencies  (Liu et al., 2024a; Zhu et al., 2024; He et al., 2024). However, their reliance on LLM-generated content inevitably introduces noise, impacting performance. (c) LLM as GNN. This paradigm directly operates LLMs as GNNs by designing structure verbalizers to convert graph data into code-like or heuristic prompts for LLM inference  (Chen et al., 2024a; Ye et al., 2024; Wang et al., 2024a, b). Yet, they fail to capture high-order connections due to the lack of an intrinsic GNN mechanism. Overall, the current decoupled approaches rely on two-stage alignment, failing to fully exploit the strengths of both models. This limitation motivates us to propose a new paradigm where LLMs inherently function as GNNs, maximizing their synergistic potential.
@@
Graph-Text Alignment in Embedding Space. Current approaches to graph-text alignment primarily operate in the embedding space. One approach uses graph encoders as prefixes, mapping graph-aware embeddings to language-based embeddings for LLMs fine-tuning in a shared space  (Huang et al., 2023, 2024). Other works adopt a two-tower architecture, leveraging contrastive learning  (Li et al., 2023a; Brannon et al., 2023; Tang et al., 2024; Jin et al., 2024b), iterative training  (Zhao et al., 2023a; Zhu et al., 2024), or knowledge distillation  (Mavromatis et al., 2023) to align distinct representations. However, these methods encounter a persistent modality gap, limiting transferability and scalability across graphs. Since TAGs inherently contain textual information, we advocate for shifting graph-text alignment to the text space.
@@
Graphs Foundation Models. A GFM aims to achieve transferability across different datasets and tasks, where the key challenge lies in finding a graph vocabulary that identifies transferable units to encode invariance on graphs  (Liu et al., 2023; Mao et al., 2024). Previous works rely on domain-specific vocabularies: GraphGPT  (Tang et al., 2024) assumes unique IDs to nodes and creates a dataset-specific vocabulary, while MoleBERT  (Xia et al., 2023) defines a molecular graph vocabulary by converting atomic properties into chemically meaningful codes. Despite their success, they lack in-context learning and cross-domain transferability. Recently, while some studies have investigated understanding and inferring graphs in natural language  (Fatemi et al., 2024; Wang et al., 2023; Zhao et al., 2023b; Liu et al., 2024a; Zhang et al., 2024b, a), none of them have attempted to establish a language-based graph vocabulary. To address this, we introduce an expressive graph vocabulary to exploit the inherent transferability of natural language, advancing the development of a versatile GFM.
@@ 3 Preliminaries
Text-Attributed Graphs. A TAG is formally represented as ${G}=\left(V,E,X\right)$, where $V$ is the set of nodes, $E$ is the set of edges. In this work, each node $v_{i}\in{V}$ is associated with a textual description ${X}_{i}=\left(x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{{n_{i}}}\right)$, where each $x_{i}^{k}\in\mathcal{X},k=1,\dots,{n_{i}}$. Here, $X$ denotes the textual attributes for nodes, and $\mathcal{X}$ represents the natural language token dictionary.
Graph Neural Networks. GNNs have emerged as state-of-the-art models in graph machine learning, predominately relying on the message-passing paradigm. In practice, a GNN first selects neighboring nodes to a target node, aggregates their representations to capture local structure, and then updates the target node’s representation. Mathematically, for a given node $v_{i}$, the $l$-th layer of a general GNN can be formulated as:
@@
where $\mathbf{h}_{i}^{(l)}$ is the node embedding of $v_{i}$ in the $l$-th layer. $\mathcal{N}_{i}$ denotes full neighbors and $\mathcal{N}_{i}^{(l)}$ is sampled neighbors in the $l$-th layer. To capture high-order connections, we stack $L$ layers and derive final embeddings as $\mathbf{h}_{i}=f_{\text{pooling}}\left(\mathbf{h}_{i}^{(1)},...,\mathbf{h}_{i}^{(L)}\right)$  (Grattarola et al., 2024). For optimization, a contrastive loss with negative sampling is commonly used in unsupervised graph learning  (Hamilton et al., 2017; Velickovic et al., 2019), expressed as:
@@ 4 Methodology
This section describes the pipeline of our proposed PromptGFM, as illustrated in Figure 3.
@@ 4.1 Graph Understanding Module
The graph understanding module aims to generate expressive representations for each node within the graph, supporting the subsequent inference module. The main challenge lies in effectively aligning the semantic and structural information, where LLMs excel in textual understanding and GNNs in structural modeling  (Li et al., 2023b; Ren et al., 2024). To bridge the gap in GNN-LLM integration, we propose prompt-based GNNs that operate LLMs as GNNs by prompting LLMs to replicate general GNN workflow within the text space.
GNN Replication with LLMs. Our priority is to design appropriate prompts that enable LLMs to function as GNNs. This requires considering three essential factors: (a) Graph Representations: How can we effectively describe the node semantics and graph structure to LLMs? (b) Graph Structure: How can we incorporate message passing to encode structural dependencies? (c) Graph Semantics: How can we distill core semantics into concise yet expressive textual representations?
As illustrated in Figure 2, we perform a fine-grained replication of GNN, i.e. prompt-based GNN. First, we initialize node representations by summarizing textual attributes, akin to look-up layers in traditional GNNs. Similarly, we sample its one-hop neighbor information to reduce computational overhead as follows:
@@
where $X^{(l-1)}=\left\{X_{0}^{(l-1)},X_{1}^{(l-1)},\dots,X_{|V|-1}^{(l-1)}\right\}$ denotes textual representations of all nodes in the previous layer, and $\left\{X_{j}^{(l-1)}\right\}$ corresponds to the selected neighbors in $\mathcal{N}_{i}$. In practice, we limit neighbor sampling to a maximum of 20 to mitigate prompt length constraints. With this context, we directly prompt LLMs to perform essential message aggregation-update process, formulated as:
$X_{i}^{(l)}$ $\leftarrow\text{Prompt}_{\text{agg-upd}}\left(\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\in\mathcal{N}_{i}^{(l)}\right\},X_{i}^{(l-1)}\right),$ (4)
where we seek flexible message passing mechanism without specification, moving beyond traditional operators like mean and weighted aggregators. In terms of optimization, we adopt an unsupervised graph learning with contrastive loss to accommodate various downstream tasks, increasing neighbor similarity while separating distant nodes. Since negative sampling is redundant in the situation, we use LLM prompts to intuitively steer this process. Prompts can be found in Appendix D.
Figure 4: An instance of graph inference module in link prediction, where language-based IDs are indexed from the graph vocabulary to generate readable instructions using task-oriented templates. We adopt multi-instruction fine-tuning framework to unify diverse graphs and tasks and learn transferable knowledge for GFMs.
After repeating $L$ rounds, we obtain the final textual representation as $T_{i}=X_{i}^{(L)}$. These representations encapsulate both semantic and structural information, effectively solving the outlined issues: (a) Graph Representations: One-hop neighbor descriptions are equivalent to an adjacency matrix, representing the entire graph structure. (b) Graph Structure: To capture high-order relationships, we iteratively invoke LLMs using the same prompts, feeding each round’s output into the next. (c) Graph Semantics: Simultaneously, we explicitly instruct LLMs to produce concise yet expressive textual representations for each node, gradually refining them for denser and richer semantics.

### v2

@@ Abstract
Text-Attributed Graphs (TAGs), where each node is associated with text descriptions, are ubiquitous in real-world scenarios. They typically exhibit distinctive structure and domain-specific knowledge, motivating the development of Graph Foundation Models (GFM) that generalize across diverse graphs and tasks. Despite large efforts to integrate Large Language Models (LLMs) and Graph Neural Networks (GNNs) for TAGs, most methods suffer from decoupled architectures with two-stage alignment, limiting their synergistic potential. Even worse, they assign out-of-vocabulary (OOV) tokens to graph nodes, leading to graph-specific semantics, token explosion, and incompatibility with task-oriented prompt templates, which hinders cross-graph and cross-task transferability. To address these challenges, we propose PromptGFM, a versatile GFM for TAGs grounded in graph vocabulary learning. PromptGFM comprises two key components: (a) Graph Understanding Module, which explicitly prompts LLMs to replicate the finest GNN workflow within the text space, facilitating seamless GNN-LLM integration and elegant graph-text alignment; (b) Graph Inference Module, which establishes a language-based graph vocabulary ensuring expressiveness, transferability, and scalability, enabling readable instructions for joint LLM fine-tuning. Extensive experiments demonstrate our superiority and transferability across diverse graphs and tasks. The code is available at https://github.com/agiresearch/PromptGFM.
@@ 1 Introduction
Graphs, characterized by their non-Euclidean structures and rich domain-specific knowledge, serve as fundamental representations of complex relational data. A wide range of real-world graphs further incorporate node-level textual information, forming Text-Attributed Graphs (TAGs), such as citation networks  [11, 18, 2], social networks  [22, 35], and molecular graphs  [52, 21]. However, existing solutions heavily rely on task- or dataset-specific training and deployment, hindering their transferability. To address this limitation, we seek to develop a Graph Foundation Model (GFM) capable of generalizing across different graphs and tasks for TAGs  [30, 56].
@@
In TAGs, Graph Neural Networks (GNNs) and Large Language Models (LLMs) are employed to process different modalities, as illustrated in Figure 1. (a) GNN for LLM. GNNs generate structure-aware node embeddings that complement original textual embeddings, improving LLM inference  [44, 5, 29]. (b) LLM for GNN. LLMs extract additional semantic features or labels from textual data, serving as supervision signals for GNN training  [8, 27, 62]. However, existing two-stage alignment pipeline, where GNNs and LLMs are only loosely coupled, struggles to fully exploit their complementary strengths, often leading to suboptimal graph-text alignment.
@@
Recently, a noteworthy trend has emerged toward using LLM as GNN, where graph verbalizers convert graph data into code-like or heuristic prompts, enabling LLMs to understand graph semantics and structure  [58, 49, 6]. However, we argue that these approaches do not qualify as true GNNs, as they omit the core message-passing mechanism. As illustrated in Figure 2(a), a traditional GNN layer includes neighbor sampling, aggregation-update, and optimization  [23, 46]. By stacking such layers, structure-less embeddings gradually evolve into structure-rich embeddings that encode higher-order dependencies. Without these essential components, a key challenge arises: Can we leverage LLMs to faithfully replicate GNNs to jointly capture both graph semantics and structure?
@@
Meanwhile, existing works intuitively treat each graph node as an out-of-vocabulary (OOV) token, resulting in graph-specific semantics and uncontrolled vocabulary explosion  [44, 58]. Worse still, due to vocabulary mismatches, ID-based node embeddings and language-based token embeddings reside in different feature spaces, causing semantic misalignment during LLM inference. This incompatibility undermines both the transferability and scalability of graph-specific knowledge across graphs and tasks. To enable knowledge transfer, an urgent question emerges: Can we replace OOV tokens with compatible and universal node representations to build a versatile GFM?
@@
A versatile GFM should be grounded in graph vocabulary learning  [3, 30, 50] with the following desired properties: (1) Expressiveness. It encapsulates both semantic and structural information across diverse graphs. (2) Transferability. Every node in any graph should be representable by a composition of fundamental units within the vocabulary. (3) Scalability. The vocabulary should be inclusive to accommodate unseen nodes, even beyond existing graphs. Since natural language is inherently an expressive medium built upon meaningful and transferable tokens  [39, 38, 37, 50], we are inspired to establish a universal graph vocabulary in the text space for node representations. Therefore, we propose PromptGFM as a GFM for TAGs as follows.
Graph Understanding Module. To function LLMs as GNNs, we initialize node features with textual attributes and prompt LLMs to explicitly replicate the fine-grained GNN workflow in text space. As shown in Figure 2, we encode local structure via one-hop neighbor descriptions and design prompts that guide a flexible aggregation-update mechanism. To support optimization, we incorporate heuristic prompts that reflect contrastive loss at each layer. Finally, iterative LLM calls simulate the message passing over the whole graph, progressively refining verbose textual features into concise yet informative textual representations. Unlike traditional GNMs that output numerical embeddings, our prompt-based GNN produces structured text encodings that preserve both node semantics and high-order connections. Consequently, LLMs can operate as GNNs, and GNNs can be interpreted as LLMs, unlocking the full potential of GNN-LLM integration and empowering elegant graph-text alignment.
Graph Inference Module. Having captured semantic and structural information through prompt-based GNNs, we decouple the resulting textual representations to establish a graph vocabulary, where each node is mapped to a finite sequence of language-based tokens, essentially as language-based IDs. This vocabulary is universally interpretable, transferable across graphs, and scalable to unseen nodes, resolving semantic irrelevance and graph-text incompatibility. Afterward, language-based IDs can be incorporated to generate readable and coherent pure-language prompts. To support generalization, we adopt a multi-instruction fine-tuning framework that gathers diverse instructions from multiple graphs and tasks, facilitating cross-graph and cross-task knowledge transfer. In conclusion, this graph vocabulary eliminates incompatibility and paves the way for general GFMs.
@@ 2 Related Works
GNN-LLM Integration. LLMs have unlocked unprecedented potential for graph machine learning, driving efforts to integrate GNNs and LLMs for modeling TAGs. (a) GNN for LLM. As structure tokenizers, GNNs and graph transformers embed graph topology into textual representations, enhancing semantical understanding in LLMs [44, 5, 29]. However, effectively coordinating GNN-LLM architectures and co-training remains a challenge. (b) LLM for GNN. LLMs assist GNNs by generating node-level and edge-level labels to address data sparsity issues  [8, 15, 56]. Meanwhile, the other line directs LLMs to produce additional features or explanations to overcome semantic deficiencies  [27, 62, 17]. However, their reliance on LLM-generated content inevitably introduces noise, impacting performance. (c) LLM as GNN. This paradigm directly operates LLMs as GNNs by designing structure verbalizers to convert graph data into code-like or heuristic prompts for LLM inference  [6, 58, 49]. Yet, they fail to capture high-order connections due to the lack of an intrinsic GNN mechanism. Overall, the current decoupled approaches rely on two-stage alignment, failing to fully exploit the strengths of both models. This limitation motivates us to propose a new paradigm where LLMs inherently function as GNNs, maximizing their synergistic potential.
@@
Graph-Text Alignment in Embedding Space. Current approaches to graph-text alignment primarily operate in the embedding space. One approach uses graph encoders as prefixes, mapping graph-aware embeddings to language-based embeddings for LLMs fine-tuning in a shared space  [19, 20]. Other works adopt a two-tower architecture, leveraging contrastive learning  [25, 1, 44], iterative training  [60, 62], or knowledge distillation  [31] to align distinct representations. However, these methods encounter a persistent modality gap, limiting transferability and scalability across graphs. Since TAGs contain textual information, we advocate for shifting graph-text alignment to the text space.
@@
Graphs Foundation Models. A GFM aims to achieve transferability across different datasets and tasks, where the key challenge lies in finding a graph vocabulary that identifies transferable units to encode invariance on graphs  [28, 30]. Previous works rely on domain-specific vocabularies: GraphGPT  [44] assumes unique IDs to nodes and creates a dataset-specific vocabulary, while MoleBERT  [55] defines a molecular graph vocabulary by converting atomic properties into chemically meaningful codes. Despite their success, they lack in-context learning and cross-domain transferability. Recently, while some studies have investigated understanding and inferring graphs in natural language  [12, 48, 61, 27], none of them have attempted to establish a language-based graph vocabulary. To address this, we introduce an expressive graph vocabulary to exploit the inherent transferability of natural language, advancing the development of a versatile GFM.
@@ 3 Preliminaries
Text-Attributed Graphs (TAGs). A TAG is defined as ${G}=\left(V,E,X\right)$, where $V$ is the set of nodes, $E$ is the set of edges, and $X$ is the set of textual node attributes. Each node $v_{i}\in{V}$ is associated with a textual description ${X}_{i}=\left(x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{{n_{i}}}\right)$, where each $x_{i}^{k}\in\mathcal{X},k=1,\dots,{n_{i}}$. Here, $\mathcal{X}$ denotes the natural language token dictionary.
Graph Neural Networks. GNNs have emerged as state-of-the-art models in graph learning, fundamentally built upon the message-passing mechanism. In practice, a GNN selects neighboring nodes to the target node, aggregates their representations to capture local structure, and updates the target node’s representation. Mathematically, for a node $v_{i}$, the $l$-th layer of a general GNN is defined as:
@@
where $\mathbf{h}_{i}^{(l)}$ is the node embedding of $v_{i}$ at the $l$-th layer. $\mathcal{N}_{i}$ denotes full neighbors and $\mathcal{N}_{i}^{(l)}$ is sampled neighbors in the $l$-th layer. To capture high-order connections, we stack $L$ layers and derive final embeddings as $\mathbf{h}_{i}=f_{\text{pooling}}\left(\mathbf{h}_{i}^{(1)},...,\mathbf{h}_{i}^{(L)}\right)$  [14]. For optimization, a contrastive loss with negative sampling is commonly used in unsupervised graph learning  [16, 47], expressed as:
@@ 4 Methodology
This section presents the pipeline of our proposed PromptGFM, which consists of a graph understanding module and a graph inference module, as illustrated in Figure 3.
@@ 4.1 Graph Understanding Module
The graph understanding module aims to generate expressive representations for each node within the graph to support downstream inference. The main challenge lies in effectively aligning semantic and structural information, where LLMs excel in textual understanding and GNNs in structural modeling  [26, 41]. To bridge the gap in GNN-LLM integration, we propose prompt-based GNNs that operate LLMs as GNNs by prompting LLMs to replicate general GNN workflow within the text space.
GNN Replication with LLMs. Our priority is to design appropriate prompts that enable LLMs to replicate GNNs. This requires considering three essential factors: (a) Graph Descriptions. How can we effectively describe the node semantics and graph structure to LLMs? (b) Graph Structure. How can we simulate message passing to encode structural dependencies? (c) Graph Semantics. How can we distill essential semantics into concise yet informative textual representations?
As illustrated in Figure 2, we implement a fine-grained replication of a general GNN workflow, i.e. prompt-based GNN. We begin by initializing node representations by summarizing their textual attributes, analogous to the lookup layer in conventional GNNs. To reduce computational overhead, we then sample one-hop neighbor information as follows:
@@
where $X^{(l-1)}=\left\{X_{0}^{(l-1)},X_{1}^{(l-1)},\dots,X_{|V|-1}^{(l-1)}\right\}$ denotes textual representations of all nodes from the previous layer, and $\left\{X_{j}^{(l-1)}\right\}$ corresponds to the selected neighbors $\mathcal{N}_{i}$. Then, we directly prompt LLMs to perform the core message aggregation-update process for each node, formulated as:
$X_{i}^{(l)}$ $\leftarrow\text{Prompt}_{\text{agg-update}}\left(\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\in\mathcal{N}_{i}^{(l)}\right\},X_{i}^{(l-1)}\right),$ (4)
where we seek flexible message passing mechanism without relying on predefined operators such as mean and weighted aggregators. A discussion on handling prompt length constraints is provided in Appendix F. In terms of optimization, we refer to unsupervised graph learning with contrastive loss to learn task-agnostic representations, which encourages similarity among neighboring nodes while pushing apart distant ones. Instead of traditional negative sampling, which becomes redundant in our setting, we leverage LLM prompts to intuitively steer the contrastive objective. Details of prompt construction are provided in Appendix D.
After repeating $L$ rounds, we obtain the final textual representation as $T_{i}=X_{i}^{(L)}$. These representations encapsulate both semantic and structural information, effectively solving the outlined issues: (a) Graph Description. One-hop neighbor descriptions are equivalent to an adjacency matrix, representing the entire graph structure. (b) Graph Structure. To capture high-order relationships, we iteratively invoke LLMs with the same prompts, feeding each round’s output into the next. (c) Graph Semantics. Simultaneously, we explicitly instruct LLMs to produce concise yet expressive textual representations for each node, gradually refining them for denser and richer semantics.